OpenStudy (monica0597):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -x - 5. y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (anonymous):

Are you familiar with the slope intercept form of a line?\[y=mx+c\]

OpenStudy (monica0597):

you mean +b? yes

OpenStudy (anonymous):

none of the above. did you write it correctly?

OpenStudy (monica0597):

y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (monica0597):

those are the choices

OpenStudy (anonymous):

what about the perpendicular line. is it correctly written too?

OpenStudy (monica0597):

its y = -1/4x - 5.

OpenStudy (anonymous):

there we go.

OpenStudy (monica0597):

sorry bout that

OpenStudy (anonymous):

it is ok

OpenStudy (anonymous):

perpendicular lines have reciprocal(fliped) slopes and have opposite signs

OpenStudy (anonymous):

when ever we have an equation of this form, and we need the line perpendicular to it. there is property obtained from Tan(A-B) for the lines, it tells us that \[m_1\times m_2= -1\]For example if you need the line perpendicular to y = 3x+5, then \[m\times 3 = -1\]this gives us \[m = \frac{-1}{3}\]this is the slope of the line perpendicular to y = 3x+5, now if we have a point ( as we do in the question) through which this perpendicular line passes then we plug it into\[ y = \frac{-1}{3}\times x+ C\]and solve for C. This will give you the equation of the perpendicular line upon substitution and simplification.

OpenStudy (anonymous):